Problem 4.6 A chemical engineer is on contract working in a relatively high-inflation country in Central America. She wishes to calculate a project's present worth with estimated costs of $30,000 now and $8000 per year for 5 year beginning one year from now. Use a real interest rate of 10% per year to make the calculations,

a. without an adjustment for inflation (10 points) and

b. considering inflation at a rate of 8% per year. (10 points)
Answer #1

a. nominal interest rate = Real interest rate + inflation + inflation * real interest rate

Here Nominal interest rate = real interest rate = 10%

NPW of the project = -30000 - 8000 * (P/A, 10%,5) = -30000 - 8000 * 3.790787 = -60326.30

b. Nominal interest rate = 0.1 + 0.08 + 0.1*0.08 = 0.188 ~ 18.8%

NPW of the project = -30000 - 8000 * (P/A, 18.8%,5)

= -30000 - 8000 * ((1 + 0.188)^5-1)/(0.188*(1 + 0.188)^5)

= -30000 - 8000 * ((1.188)^5-1)/(0.188*(1.188)^5)

= -30000 - 8000 * 3.071337

= -54570.70